if the perimeter of a square is 48 yards what is the are of the square

Answer by Mr. Math Wiz(7)

-Square has four sides so you would divide it by 4
-48/4=12
-To find area, you square the side or multiply it by itself
-12^2 or 12*12=144
-The answer is 144